OpenStudy (anonymous):

An expression is shown below: f(x) = -16x^2 + 22x + 3 Part A ~ What are the x-interecpts of the graph f(x)? Show your work. Part B ~ Is the vertex of the graph f(x) going to be a maximum or a minimum? what are the coordinates of the vertex? Justify your answer and show your work. Part C ~ What are the steps you would use to graph f(x)? Make sure you can answer the obtained in Part A and Part B to draw the graph. It's an essay question, so you're not really making a graph. It would be if you had to draw it.

OpenStudy (solomonzelman):

part A. f(x) = -16x^2 + 22x + 3 0 = -16x^2 + 22x + 3 can you solve this quadratic equation ?

OpenStudy (anonymous):

No, i'm not sure how to solve quadric equations. It's my unit 4, which is the next unit i'm going to be working on.

OpenStudy (solomonzelman):

multiply the equation times -1, you get... 16x^2 - 22x - 3 = 0 use the quadratic formula.

OpenStudy (anonymous):

So you'd get... -16x^2 + 22 + 3 =0 Right?

OpenStudy (solomonzelman):

I got that when I set y (or "f(x)" - same thing) equal zero. I multiplied everything times -1 and now, can you solve for x using a quadratic formula ?

OpenStudy (anonymous):

Do I just add square 16, subtract 22 and 3 and divide them? I don't know what a quadric formula is. I haven't learned anything about it yet.

OpenStudy (solomonzelman):